# Training Language Models on Synthetic Edit Sequences Improves Code Synthesis

### v2

@@ 4 Related Work
Finetuning on Edits
Several works have investigated finetuning code LLMs on edit data. Notably, Muennighoff et al. (2023) instruction tune models on a 4TB dataset of GitHub commits pairing code changes with human instructions. Relatedly, Li et al. (2024) use GitHub commit data sourced from Python repositories to generate code editing instruction data with GPT 3.5/ChatGPT. Both of these works specifically focus on better-equipping LLMs for natural language-prompted code editing tasks, in which a model is explicitly prompted to generate an edit in response to a natural language specification. Our work differs in three important ways: first, we study edit sequences rather than single edits; second, we train LLMs to predict edits implicitly during code synthesis; third, our synthetic edit generation algorithm does not rely on the existence of any kind of commit data.
@@
Inference-Time Compute Scaling
@@ 5 Discussion, Limitations, and Conclusion
This paper introduces an algorithm, LintSeq, for generating synthetic code edit sequences from existing programs. LintSeq enables LLM reasoning settings like code synthesis to be re-parameterized at the data-level as sequential edit generation tasks. The algorithm is parameter-free, requires only CPU to run, and makes no assumptions about the content or structure of code files.
Re-parameterizing code generation with edits has a few immediate benefits. For example, it makes code generation with LLMs much more controllable at the prompt-level (Appendix A.3) and it reduces the cost of predicting useful and correct code insertions with models, since synthetic edit-trained LLMs do not need to be prompted to re-synthesize entire programs from scratch (Section 2.4).
@@
    Tiny LMs can be efficiently finetuned to synthesize Python programs with edit sequences via LintSeq data. This results in state-of-the-art code benchmark performance for models that can be trained “on device” (Sections 3.1 and 3.3.1).
@@
    Across other tested models from the Phi, Gemma, and Llama families, finetuning on LintSeq data also improves the diversity of zero-shot generations, boosting the inference-time scaling of coverage on HumanEval and MBPP at fixed sample counts (Section 3.3.2).
@@
    On HumanEval, the cumulative inference cost of repeatedly sampling from small edit sequence LLMs is similar to sampling once from larger LLMs and yields coverage that is competitive with GPT-4, GPT-4-Omni, and Llama 3.1 405B (Figure 1, Appendix F.4).
@@
    Ablating linter-guidance from LintSeq hurts the quality, diversity, and correctness of code synthesized by instruction finetuned models (Section 3.4).
@@
First, as currently formulated, LintSeq can only be used to generate synthetic sequences of insertion edits. This is a consequence of the parameter-free nature of the algorithm – every edit in a LintSeq sequence reflects an existing line of code in the source file used to generate it. As a result, LintSeq edit data can only be used to train models for synthesis-like tasks, rather than for refinement.
Second, our experiments with LintSeq study code synthesis in Python only. We look forward to testing LintSeq with other programming languages, verifiers, and problems in future work. One other exciting setting where LintSeq might improve the tradeoff between model generation quality and inference-time compute is in mathematical reasoning and formal theorem proving.
@@ Acknowledgements
Appendix A More on Edit Sequences and Diffs
A.1 Reading Unix Diffs
We provide a guide to reading Unix-style diffs below in Figure 7. The diff shown in this figure is computed using the Python library difflib, which is the implementation that we use to compactly represent edits in our synthetic data generation experiments. Note that the total extra tokens present in an insertion edit sequence representation of a program scales with the number of program lines $L$, and can be upper-bounded as $T_{\text{diff}}\leq L\cdot((\text{chars in ``decorator''})+(\text{extra chars per line in ``body''}))=16L$.
@@ A.1 Reading Unix Diffs
Figure 7: The anatomy of a Unix diff: A diagrammatic visualization of the different parts of a Unix-style diff, as computed by difflib. The body of a diff can consist of multiple line deletions, followed by multiple line insertions. The decorator portion of the diff respectively indicates the location and size of these deletions and insertions, if any. Like the diff shown above, the edits in synthetic edit sequences generated by LintSeq consist of line insertions only.
A.2 Resolving Edit Sequences
@@
During inference, LMs that have been finetuned on LintSeq instruct data will synthesize code via edit sequences, outputting text strings that consist of a sequence of consecutive Python diffs interleaved with newline characters and “<|diff|>” tokens, similar to  Piterbarg et al. (2024). Each of these diffs will be structured as shown in Figure 7, if correctly formatted by the language model.
@@
Figure 1 shows a code edit sequence generation from a LintSeq instruction finetuned LM and the corresponding resolved, executable Python program.
A.3 Controllability of Code Synthesis with Edit Sequence LMs
@@
The structure of Unix-style diffs affects the downstream controllability of code synthesis with models that have been trained on edit sequence re-parameterized programs. As shown in Figure 7, the first line of every diff is a decorator that describes the location and the numbers of lines changed by the edit. During inference, autoregressive language models that have been trained on Unix-style diffs with this format can be prompted to predict an edit in any desired target location within the program being synthesized by “intervening” on a model generation.
A.4 Future Work: Searching in Edit Space
@@
If we apply the lens of reinforcement learning or search to this setting, we might say that re-parameterizing the code data used to train a language model re-parameterizes the model’s generative action space. It is possible that combining edit sequence LMs with more sophisticated decoding mechanisms, inference-time search, and/or interactive post-training may result in even larger improvements to the quality of generated code than those of the zero-shot code synthesis settings studied in this paper. We look forward to testing this hypothesis in future work.
Appendix B Evaluation
@@
HumanEval (Chen et al., 2021) and Mostly-Basic Programming Problems (MBPP) (Austin et al., 2021) are two of the most studied benchmarks for evaluating code LMs (Liu et al., 2023). These benchmarks probe the code synthesis capabilities of models, and consist of pairs of natural language program descriptions and test-cases. We employ the extended MBPP test cases released as MBPP(+) by Liu et al. (2023) to add additional rigour to our testing procedure. All of the code LMs that we compare our models against evaluate HumanEval performance using the original set of benchmark test cases; for consistency, we employ these same test cases in our evaluations when comparing the performance of our models to the reported scores of external LMs.
@@
During testing on both HumanEval and MBPP(+), LMs are prompted to generate outputs using the natural language descriptions of target programs. Their outputs are then evaluated on the paired test cases. A generation is considered “correct” if and only if it passes all of the test cases upon execution, subject to a fixed timeout setting. Previous works on code synthesis with language models report scores on HumanEval and MBPP(+) across samples. The most common of these metrics is known as pass@k (Chen et al., 2021; Austin et al., 2021; Li et al., 2022b; Wang et al., 2023b). This is the metric that we use to report and compare model performance throughout this paper. Our implementation of HumanEval and MBPP(+) evaluations mimics the Big Code Evaluation Harness by Ben Allal et al. (2022). We do not allow models to use chain-of-thought during generation.
B.1 Prompting

### v3

@@ 4 Related Work
Training on Edits
Many works have studied edit generation with language models. Yin et al. (2018) cast the edit representation problem as an autoencoding task and show that neural network models can learn to capture the structure and semantics of edits, while Gu et al. (2019) introduce a partially autoregressive model for generating insertion and deletion edits that is trained with adversarial imitation learning. Guo et al. (2021) use reinforcement learning to train LMs to generate code with “holes” that represent high uncertainty tokens, and to edit the contents of these “holes” later on.
More recently, several works have investigated finetuning off-the-shelf pre-trained language models on large-scale edit data. Berabi et al. (2021) use a linter to detect errors in code, and finetune a T5 model (Raffel et al., 2020) to correct code by leveraging error messages. Muennighoff et al. (2023) and Cassano et al. (2023) instruction tune models on datasets of GitHub commits pairing code changes with human instructions. Relatedly, Li et al. (2024) use GitHub commit data sourced from Python repositories to generate code editing instruction data with GPT 3.5/ChatGPT. All of these works specifically focus on better-equipping LMs for natural language-prompted code editing tasks, in which a model is explicitly prompted to generate an edit in response to an error message or a natural language specification. Our work differs in three important ways: first, we study edit sequences rather than single edits; second, we train LMs to predict edits implicitly during code synthesis; third, our synthetic edit generation algorithm does not rely on the existence of any kind of commit data.
@@
Scaling Up Test-Time Compute
@@ 5 Discussion, Limitations, and Conclusion
This paper introduces an algorithm, LintSeq, for generating synthetic code edit sequences from existing programs. LintSeq enables code synthesis to be re-parameterized at the data-level as sequential edit generation tasks. The algorithm is parameter-free, requires only CPU to run, and makes no assumptions about the content or structure of source code files.
Re-parameterizing code generation with edits has a few immediate benefits. For example, it makes code generation with LMs more controllable at the prompt-level (Appendix B.3) and it reduces the cost of predicting useful and syntactically correct code insertions with models, since synthetic edit-trained LMs do not need to be prompted to re-generate full programs from scratch (Section 2.5).
@@
    Tiny LMs pre-trained for code understanding can be efficiently fine-tuned to synthesize programs edit-by-edit via LintSeq data. This results in competitive performance on HumanEval and MBPP(+) compared to existing code LMs of similar scale (Sections 3.1 and 3.3.1).
@@
    On larger models from the Phi 3, Gemma 2, and Llama 3.1 families that were pretrained for general natural language understanding, tuning on LintSeq data either improves or preserves the quality of pass@1 generations compared to standard tuning (Section 3.3.2).
@@
    LintSeq also improves test-time compute scaling laws for code synthesis on instruction fine-tuned Phi 3, Gemma 2, and Llama 3.1 models, suggesting that edit sequence LMs consistently generate more meaningfully diverse programs compared to baselines, even on challenging benchmarks like CodeContests (Section 3.3.2).
@@
    Ablating the linter from LintSeq hurts the quality and syntactical correctness of code synthesized by edit sequence TinyCodeLMs. This suggests that the structured nature of edits sampled with LintSeq is important for downstream LM performance (Section 3.4).
@@
First, as currently formulated, LintSeq can only be used to generate synthetic sequences of insertion edits. This is a consequence of the parameter-free nature of the algorithm – every edit in a LintSeq sequence reflects an existing line of code in the source file used to generate it. As a result, models that are fine-tuned exclusively on data sampled with LintSeq cannot be used for code editing tasks involving deletion edits. One simple way to circumvent this limitation might be by mixing LintSeq synthetic edit sequences with human edit data during instruction fine-tuning via datasets like CommitPackFT (Muennighoff et al., 2023), which contain examples of deletions. An alternate approach might be to follow-up supervised instruction fine-tuning on LintSeq synthetic data with reinforcement learning in order to train models to interleave insertions with deletions when necessary.
Second, the experiments that we conducted with LintSeq in this paper studied code synthesis in Python only. LintSeq can be similarly used for generating synthetic edit sequences for code written in other programming languages by swapping out the linter using during edit sampling.
Finally, we used LintSeq to refactor an instruction fine-tuning dataset in this work. However, by design, the algorithm can be run on any corpus of source code data, such as The Stack (Kocetkov et al., 2022) or The Stack-v2 (Li et al., 2023). In future work, we hope to explore using LintSeq to train LMs to write code edit-by-edit on larger, pre-training scale datasets.
@@
Reproducibility Statement
In the supplementary materials accompanying this submission, we provide a Python implementation of LintSeq as well as instructions and code supporting data generation, processing, pretraining, and fine-tuning experiments. We also provide thorough textual descriptions of all experimental procedures in the Appendix. Appendix C describes prompting and model evaluation, while Appendices D and E detail all of the hyperparameters, procedures, and open-source datasets that we employ for obtaining the results reported throughout Section 3. Finally, Appendix A.4 provides references and data for reproducing the results plotted in Figure 1.
@@ Acknowledgements
Appendix A Additional Results
A.1 Empirics of Processing Code Data with LintSeq
Figure 6: Empirics of processing code data with LintSeq. Left: Lines per example in a dataset of instruction fine-tuning data for Python synthesis before and after processing with LintSeq via the linter pylint (see Section 3.2). LintSeq processing adds lines of diff metadata to examples (see Appendix B). Right: The corresponding edit counts per synthetic code edit sequence. On a dataset of short programs (14 lines of code, on average), the mean LintSeq edit sequence contains four edits.
A.2 Comparing LintSeqInstruct to RandSeqInstruct TinyCodeLMs on HumanEval and MBPP(+)
Table 2: Edit sequence TinyCodeLM results on HumanEval at high sampling temperatures: We tune sampling parameters for edit sequence variants of TinyCodeLM over temperatures (1, 1.1, 1.2), top-p (0.95, 1.0), and min-p (0, 0.05) with $n=64$ completions per problem and report the best pass@k value obtained from each model variant. We also report standard error for each score.
HumanEval
tinycodeLM-RandSeqInstruct 150M 4.5 $\pm$ 0.4 10.3 $\pm$ 0.5 12.2 $\pm$ 0.5 14.4 $\pm$ 0.6 18.8 $\pm$ 0.6
tinycodeLM-LintSeqInstruct 150M 6.4 $\pm$ 0.5 13.9 $\pm$ 0.5 16.8 $\pm$ 0.6 19.5 $\pm$ 0.6 23.6 $\pm$ 0.6
tinycodeLM-RandSeqInstruct 400M 6.0 $\pm$ 0.4 11.7 $\pm$ 0.5 13.9 $\pm$ 0.6 16.4 $\pm$ 0.6 20.8 $\pm$ 0.6
tinycodeLM-LintSeqInstruct 400M 8.4 $\pm$ 0.4 16.6 $\pm$ 0.6 19.7 $\pm$ 0.6 22.8 $\pm$ 0.6 27.2 $\pm$ 0.6
Table 3: Edit sequence TinyCodeLM results on MBPP(+) at high sampling temperatures: As above, we tune sampling parameters for all fine-tuned TinyCodeLM variants over temperatures (1, 1.1, 1.2), top-p (0.95, 1.0), and min-p (0, 0.05) with $n=64$ completions per problem and report the best pass@k value obtained from each model variant. Standard error is indicated with “$\pm$.”
MBPP(+)
tinycodeLM-RandSeqInstruct 150M 6.5 $\pm$ 0.3 17.2 $\pm$ 0.4 22.6 $\pm$ 0.4 27.9 $\pm$ 0.5 34.4 $\pm$ 0.5
tinycodeLM-LintSeqInstruct 150M 8.6 $\pm$ 0.3 19.5 $\pm$ 0.4 24.5 $\pm$ 0.5 29.0 $\pm$ 0.5 35.1 $\pm$ 0.5
tinycodeLM-RandSeqInstruct 400M 10.2 $\pm$ 0.4 20.8 $\pm$ 0.4 25.4 $\pm$ 0.5 29.9 $\pm$ 0.5 36.2 $\pm$ 0.5
tinycodeLM-LintSeqInstruct 400M 14.7 $\pm$ 0.4 25.8 $\pm$ 0.5 29.6 $\pm$ 0.5 33.9 $\pm$ 0.5 39.7 $\pm$ 0.5
A.3 HumanEval, MBPP(+), CodeContests, DS-1000, and BigCodeBench results for LintSeq vs Baseline Instruction Tuned Gemma 2, Phi-3, and Llama 3.1 Models
Table 4: Gemma 2, Phi-3, and Llama 3.1 results on HumanEval at high sampling temperatures. We report the best pass@k value obtained from each model variant at high sampling temperatures, sweeping over temperature values (1, 1.1, 1.2), top-p (0.95, 1.0), and min-p (0, 0.05). We generate $n=64$ completions per problem and report standard error for each estimated score.
HumanEval
Model Variant Size pass@1 pass@5 pass@10 pass@20 pass@50
Gemma-2-Instruct 2.6B 15.3 $\pm$ 0.6 22.0 $\pm$ 0.6 25.2 $\pm$ 0.6 31.6 $\pm$ 0.6 41.7 $\pm$ 0.7
Gemma-2-LintSeqInstruct 2.6B 22.0 $\pm$ 0.6 34.8 $\pm$ 0.6 41.4 $\pm$ 0.6 48.2 $\pm$ 0.7 55.5 $\pm$ 0.7
Phi-3-Mini-Instruct 3.8B 35.2 $\pm$ 0.6 49.7 $\pm$ 0.6 55.1 $\pm$ 0.7 59.2 $\pm$ 0.7 62.2 $\pm$ 0.7
Phi-3-Mini-LintSeqInstruct 3.8B 38.4 $\pm$ 0.6 63.3 $\pm$ 0.6 72.4 $\pm$ 0.6 79.9 $\pm$ 0.6 87.3 $\pm$ 0.5
Llama-3.1-Instruct 8B 38.4 $\pm$ 0.6 51.3 $\pm$ 0.7 56.2 $\pm$ 0.7 60.2 $\pm$ 0.7 64.2 $\pm$ 0.7
Llama-3.1-LintSeqInstruct 8B 38.5 $\pm$ 0.6 62.2 $\pm$ 1.6 72.6 $\pm$ 1.6 75.7 $\pm$ 0.6 82.7 $\pm$ 0.6
Phi-3-Med-Instruct 14B 50.2 $\pm$ 0.6 68.4 $\pm$ 0.6 73.5 $\pm$ 0.6 77.3 $\pm$ 0.6 81.4 $\pm$ 0.6
Phi-3-Med-LintSeqInstruct 14B 49.7 $\pm$ 0.6 75.0 $\pm$ 0.6 81.6 $\pm$ 0.6 85.9 $\pm$ 0.6 89.6 $\pm$ 0.5
Table 5: Gemma 2, Phi-3, and Llama 3.1 results on MBPP(+) at high sampling temperatures. Exactly as above, we sweep over temperature (1, 1.1, 1.2), top-p (0.95, 1.0), and min-p (0, 0.05) and report the best pass@k value obtained from each model variant. We generate $n=64$ completions per problem and report standard error for each estimated score.
MBPP(+)
Model Variant Size pass@1 pass@5 pass@10 pass@20 pass@50
Gemma-2-Instruct 2.6B 20.5 $\pm$ 0.4 30.8 $\pm$ 0.5 34.3 $\pm$ 0.5 37.6 $\pm$ 0.5 41.6 $\pm$ 0.5
Gemma-2-LintSeqInstruct 2.6B 28.2 $\pm$ 0.5 40.1 $\pm$ 0.5 44.5 $\pm$ 0.5 48.6 $\pm$ 0.5 52.8 $\pm$ 0.5
Phi-3-Mini-Instruct 3.8B 31.9 $\pm$ 0.5 42.5 $\pm$ 0.5 46.3 $\pm$ 0.5 49.8 $\pm$ 0.5 53.6 $\pm$ 0.5
Phi-3-Mini-LintSeqInstruct 3.8B 37.2 $\pm$ 0.5 51.4 $\pm$ 0.5 56.1 $\pm$ 0.5 60.3 $\pm$ 0.5 66.0 $\pm$ 0.5
Llama-3.1-Instruct 8B 37.4 $\pm$ 0.5 50.2 $\pm$ 0.5 53.6 $\pm$ 0.5 56.6 $\pm$ 0.5 60.0 $\pm$ 0.5
Llama-3.1-LintSeqInstruct 8B 40.3 $\pm$ 0.5 56.2 $\pm$ 0.5 61.1 $\pm$ 0.5 65.5 $\pm$ 0.5 69.4 $\pm$ 0.5
Phi-3-Med-Instruct 14B 37.7 $\pm$ 0.5 50.4 $\pm$ 0.5 54.0 $\pm$ 0.5 57.0 $\pm$ 0.5 60.1 $\pm$ 0.5
Phi-3-Med-LintSeqInstruct 14B 39.1 $\pm$ 0.5 55.2 $\pm$ 0.5 60.7 $\pm$ 0.5 65.4 $\pm$ 0.5 71.1 $\pm$ 0.5
Table 6: Gemma 2, Phi-3, and Llama 3.1 results on CodeContests. We sweep over temperature (0.5, 0.6) and use top-p $=1$, min-p $=0$, and $n=128$, and report the best pass@k value obtained from each model variant in the table below. We also report standard error for each estimated score.
CodeContests
Model Variant Size pass@1 pass@50 pass@100
Gemma-2-Instruct 2.6B 0.05 $\pm$ 0.05 1.56 $\pm$ 0.26 2.26 $\pm$ 0.30
Gemma-2-LintSeqInstruct 2.6B 0.61 $\pm$ 0.16 5.71 $\pm$ 0.37 7.03 $\pm$ 0.40
Phi-3-Mini-Instruct 3.8B 1.80 $\pm$ 0.22 14.86 $\pm$ 0.45 18.59 $\pm$ 0.49
Phi-3-Mini-LintSeqInstruct 3.8B 2.76 $\pm$ 0.26 19.10 $\pm$ 0.48 22.93 $\pm$ 0.51
Llama-3.1-Instruct 8B 2.68 $\pm$ 0.28 11.21$\pm$ 0.44 12.80 $\pm$ 0.46
Llama-3.1-LintSeqInstruct 8B 2.92 $\pm$ 0.27 17.86 $\pm$ 0.47 21.82 $\pm$ 0.51
Phi-3-Med-Instruct 14B 3.22 $\pm$ 0.27 16.50 $\pm$ 0.47 19.45 $\pm$ 0.50
Phi-3-Med-LintSeqInstruct 14B 3.02 $\pm$ 0.25 19.09 $\pm$ 0.48 23.11 $\pm$ 0.51
Table 7: Gemma 2, Phi-3, and Llama 3.1 pass@1 results on DS-1000. We use the same sampling hyperparameters as Luo et al. (2023) and Wei et al. (2024b) to evaluate instruction tuned models.
Model Variant Size DS-1000, pass@1
Gemma-2-Instruct 2.6B 2.5
Gemma-2-LintSeqInstruct 2.6B 3.8
Phi-3-Mini-Instruct 3.8B 8.6
Phi-3-Mini-LintSeqInstruct 3.8B 15.5
Llama-3.1-Instruct 8B 14.5
Llama-3.1-LintSeqInstruct 8B 16.2
Phi-3-Med-Instruct 14B 21.8
Phi-3-Med-LintSeqInstruct 14B 24.2
Table 8: Gemma 2, Phi-3, and Llama 3.1 pass@1 results on BigCodeBench (Instruct). We use greedy decoding to evaluate instruction tuned models.
Model Variant Size BigCodeBench Instruct, pass@1
Gemma-2-Instruct 2.6B 5.44
Gemma-2-LintSeqInstruct 2.6B 6.32
Phi-3-Mini-Instruct 3.8B 20.79
Phi-3-Mini-LintSeqInstruct 3.8B 21.58
Llama-3.1-Instruct 8B 21.46
Llama-3.1-LintSeqInstruct 8B 20.53
Phi-3-Med-Instruct 14B 24.65
Phi-3-Med-LintSeqInstruct 14B 28.16
A.4 Computing Pass@k vs Total Test-Time FLOPs
In Figures 1(right) and 4, we plot the percentage of problems solved by any attempt (i.e. pass@k) on HumanEval, MBPP, and CodeContests as a function of total test-time FLOPs used during sampling for LintSeq vs baseline instruction fine-tuned models. Raw “pass@k” estimates are also included in Tables 4, 5, and 8, representing the best scores achieved by each model variant after tuning sampling hyperparameters.
We compute total test-time FLOPs using the approximations below, which are drawn from Kaplan et al. (2020). These approximations conservatively estimate the cumulative inference costs of synthesizing solutions to all of the problems in the test set of each benchmark. The models that we compare are all dense transformers, where the majority of the parameters are used in matrix multiplications.
We determine the quantities $T_{\text{avg-total-tokens-per-sample}}$ for each model variant at a particular “pass@k” by computing token counts over all sets of samples per problem.
Note that edit sequence (i.e. LintSeqInstruct fine-tuned) LMs have slightly higher average token counts per sample due to presence of “diff” descriptor tokens in generations (see Appendix B).
Appendix B More on Edit Sequences and Diffs
B.1 Reading Unix Diffs
[…]
@@ B.1 Reading Unix Diffs
Figure 7: The anatomy of a Unix diff: A diagrammatic visualization of the different parts of a Unix-style diff, as computed by difflib. The body of a diff can consist of multiple line deletions, followed by multiple line insertions. The decorator portion of the diff shows the location and size of these deletions and insertions, if any. Like the diff shown above, the edits in synthetic edit sequences generated by LintSeq consist of line insertions only.
B.2 Resolving Edit Sequences
@@
During inference, LMs that have been fine-tuned on LintSeq instruct data will iteratively synthesize programs by generating edits i.e., outputting text that consists of a sequence of consecutive Python diffs interleaved with newline characters and “<|diff|>” tokens, similar to  Piterbarg et al. (2024). If correctly formatted by the LM, these diffs will be structured as shown in Figure 7.
@@
Figure 1 shows a code edit sequence generation from a LintSeq instruction fine-tuned LM and the corresponding resolved, executable Python program.
B.3 Controllability of Code Synthesis with Edit Sequence LMs
@@
The structure of Unix-style diffs affects the downstream controllability of code synthesis with models that have been trained on edit sequence re-parameterized programs. As shown in Figure 7, the first line of every diff is a decorator that describes the location and the number of lines changed by the edit. During inference, autoregressive language models that have been trained on diffs with this format can be prompted to predict an edit in a target location by intervening on a model generation.
B.4 Future Work: Searching in Edit Space
@@
If we apply the lens of reinforcement learning or search to this setting, we might say that re-parameterizing the code data used to train a language model re-parameterizes the model’s action space. It is possible that combining edit sequence LMs with more sophisticated decoding mechanisms, test-time search, and/or reinforcement learning may result in even larger improvements to the quality of generated code than those of the zero-shot code synthesis settings studied in this paper. We look forward to testing this in future work.
Appendix C Evaluation
@@
HumanEval (Chen et al., 2021) and Mostly-Basic Programming Problems (MBPP) (Austin et al., 2021) are two of the most studied benchmarks for evaluating code LMs (Liu et al., 2023). These benchmarks probe the code synthesis capabilities of models, and consist of pairs of natural language program descriptions and test-cases. We employ the extended MBPP test cases released as MBPP(+) by Liu et al. (2023) to add additional rigour to our testing procedure. The code LMs that we compare our TinyCodeLM models against in Table 1 evaluate HumanEval performance using the original set of benchmark test cases; for consistency, we employ these same test cases in all of our evaluations.
Our evaluations on the harder benchmarks CodeContests, DS-1000, and BigCodeBench(Instruct) use exactly the same sets of problem descriptions and test cases as those introduced by Li et al. (2022b), Lai et al. (2023), and Zhuo et al. (2024).
@@
During testing on each benchmarks, LMs are prompted to generate outputs using the natural language descriptions of target programs. Their outputs are then evaluated on the paired test cases. A generation is considered “correct” if and only if it passes all of the test cases upon execution, subject to a fixed timeout setting. Previous works on code synthesis with language models report scores across samples. The most common of these metrics is known as pass@k (Chen et al., 2021; Austin et al., 2021; Li et al., 2022b; Wang et al., 2023b). This is the metric that we use to report and compare model performance throughout this paper.
C.1 Prompting